# UDA: A Benchmark Suite for Retrieval Augmented Generation in Real-world Document Analysis

### v1

@@ 1 Introduction
The initial findings from our experiments can be pivotal but limited. We are actively updating the benchmark suite and will incorporate more state-of-the-art RAG solutions and LLMs in our benchmark. We hope such efforts will shed light on future research and production of RAG- and LLM-based document analysis.
@@ 3 Dataset: UDA
In this section, we outline the construction of UDA. Each data item within UDA is logically structured as a triplet $(D,q,a)$, where $D$ represents a complete unstructured document, $q$ denotes a question raised from the document, and $a$ signifies the ground truth answer (refer to the data example in Appendix  A). To mirror the authenticity of real-world applications, the documents are retained in their original file formats without parsing or segmentation.
@@
Arithmetic What was the percentage increase in cash dividend from 2015 to 2016? $(0.29-0.25)\div 0.25*100\%=16\%$
Data Sources
@@
We first collect the Q&A labels from the open-released datasets (i.e., source datasets), which are all annotated by human participants. Specifically, our FinHybrid is based on the financial numerical reasoning dataset FINQA [11], which is constructed based on the public earnings reports of S&P 500 companies. TatHybrid is derived from TAT-DQA [63], whose Q&A pairs are accompanied by the document snapshot of 1 to 3 pages from public financial annual reports. Both PaperTab and PaperText are based on Qasper [12], a reading comprehension dataset based on NLP research papers. FetaTab is built upon FetaQA [34], a question-answering dataset for tables from Wikipedia pages.
NqText is derived from the widely used Q&A dataset, Natural-Questions [27], which uses the Wikipedia pages as context. Its questions are collected from the Google Search engine, and the answers are human-annotated.
@@
We conduct a series of essential constructing actions after collecting the Q&A labels from the source datasets. The integrity of original documents is crucial for the fidelity of document analysis. However, most of the source datasets only offer well-parsed and segmented partial content without the complete document.
To address this problem, we perform a comprehensive source-document identification process. We meticulously search for, verify, and collect the complete original document files, according to the metadata or content fragments from every source dataset. Then we embark on a rigorous matching and reorganization effort, forming complete triplet data pairs, i.e., document-question-answer. Additionally, we categorize queries by the source of factual evidence, filter out Q&As without available answers, convert token-based data patterns to natural language, unify data formats and structures across datasets, and design specific LLM-prompts for each dataset after experimental attempts.
@@ 4 Benchmarks and Evaluations
To evaluate the quality of LLM-generated answers, we apply widely accepted span-level F1-score [40] in PaperTab, PaperText, FetaTab, and NqText datasets, where ground-truth answers are in natural language and the source datasets also utilize this metric. We treat the prediction and ground truth as bags of words and calculate the F1-score to measure their overlap. In financial analysis, the assessment becomes more intricate due to numerical values. For the TatHybrid dataset, we adopt the numeracy-focused F1-score, introduced by Zhu, et al. [64], which considers the scale and the plus-minus of numerical values. In the FinHybrid dataset, where answers are always numerical or binary, we rely on the Exact-Match metric but allow for a numerical tolerance of $1\%$, accounting for rounding discrepancies. To assess the effectiveness of retrieval strategies, we identify the factual evidence in retrieved chunks using the relative length of the Longest Common Subsequence (LCS) [38] instead of the exact match, because extracted PDF data chunks often include extraneous symbols that can hinder exact matches while still containing crucial evidence.
@@
Following prior works in [3, 57, 65], we focus on zero-shot LLM generation, yet adding an extra formatting example to align the output with the desired pattern (refer to Appendix B.1). For the GPT-4 model, we leverage the Azure-OpenAI API to access GPT4-Turbo-1106-Preview with the context window of 128k. Other open-source models are obtained from Huggingface, and we always use the instruct-tuned version. The inference is done on 4 NVIDIA-A100 GPUs. To reduce the compute costs, we randomly sample 1201 documents (in PDF format) accompanied by 2503 question-answer pairs to form our evaluation set (detailed in Table  4). We believe it serves as a practical performance indicator for real-world scenarios.
@@ 4.1 Evaluating Data Parsing
We evaluate various parsing methods to extract tabular information from PDF files and analyze their influence on the downstream tasks. We utilize the question set from PaperTab and the table-based questions from FinHybrid. As shown in Figure 2, each question is paired with a PDF page that contains the clue tables; doing so prevents inaccurate retrievals. The tabular data are parsed into text and merged with the rest of the text content as the input context to the LLM.
@@
We evaluate several existing approaches of table parsing: (1) Raw text extraction, which employs a PDF text extractor [39] to extract all the characters. (2) Classic Computer Vision (CV) based approach, which often performs layout detection and OCR extraction at the same time. We follow [53] to use Yolox [16], Tesseract [48] and TableTransformer [49] models together. (3) CV + LLM method, which further employs an LLM to transform the outputs of (2) into Markdown tables. (4) For the advanced multi-modal approach, we employ the latest GPT-4-Omni [35] to convert image-based document tables into Markdown format. (5) The FinHybrid dataset provides the verified well-parsed tables, which serve as the parsing ground truth.
Table 5: Performance scores (EM or F1) of LLMs using varying parsing strategies on table-based Q&A tasks. Due to a lack of well-parsed table data in PaperTab dataset, there are no results.
@@
PaperTab (F1) GPT-4-Turbo - 42.4 42.0 41.0 40.6
Llama-3-8B - 32.4 35.1 33.4 32.2
Table 5 reveals that GPT-4-Omni outperforms other approaches, while surprisingly, raw text extraction also yields decent results. We found that the queried tables are relatively simple; the structural markers from the raw text, such as line-breakers and space, are often adequate for LLMs to understand the table. Classic CV methods, if not meticulously tuned, may struggle in handling non-standard table presentations, i.e., edge cases (see an example in Figure  3). Additionally, employing GPT-4-Omni directly for question-answering scores 69.8 and 35.4, lower than sequentially parsing and generating with GPT-4 (i.e., 72.4 and 42.4).
@@
We also observe from the FinHybrid dataset that while the GPT-4 model shows a modest 5.7% improvement with well-parsed data, the much smaller Llama-3-8B offers a significant 15% enhancement, suggesting that compact models with a limited capability of parsing table layouts, may benefit more from enhanced parsing.
@@
Evaluations here suggest (1) CV-based parsing methods may require adaptation for edge cases before they can be useful; (2) smaller LLMs may be impacted more by uncleaned input data.
@@ 4.2 Indexing and Retrieval
We use the relative length of the Longest Common Subsequence (LCS) to demonstrate the presence of human-annotated evidence in retrieved chunks (more experimental details in Appendix B.2). As shown in Table  6, OpenAI’s text-embedding-3-large model excels in all datasets except for FinHybrid, where the simpler BM-25 approach intriguingly outperforms. This could be attributed to the fact that financial queries often contain more precise details, such as dates or keywords; this aligns well with the direct keyword-matching of BM-25.
@@
We also conduct end-to-end experiments to verify the impact of the retrieval quality. We evaluate the answer quality with top-5 chunks retrieved using OpenAI embeddings or with the human-annotated evidential chunks. As shown in Table 7, providing more relevant context to LLMs improves the answers in most cases, particularly in arithmetic-reasoning tasks such as FinHybrid and TatHybrid. Interestingly, for the FinHybrid dataset, the Llama-3-8B model achieved a score of 51.0 when given accurate context, outperforming GPT-4 with retrieved chunks. However, for knowledge-based questions from the NqText and FetaTab, the answer quality remains less affected. This is because complex numerical reasoning demands more precise evidence for accurate arithmetic operations, whereas LLMs can leverage a wide range of narrative information to derive answers to knowledge-based questions.
@@
We compare RAG-based methods with long-context LLMs, utilizing GPT-4-Turbo with a 128k context window and Qwen-1.5-7B with a 32k context window. Due to the high cost of long context inference, we conduct this experiment on a subset of 600 documents (more details in Appendix  B.3). The results are demonstrated in Table  8.
@@
Remark. We notice that for free-form or knowledge-based tasks (i.e., paper-based and wiki-based Q&A), RAG and long-context solutions demonstrate comparable capability. Conversely, in tasks with more numerical reasoning (i.e., financial Q&A), long-context LLMs fail to match the performance of RAG. In such use cases, the excess of verbose content might hinder long-context LLMs in pinpointing facts and performing numerical reasoning effectively (see an example in Table  10).
@@ 4.4 Evaluating Chain-of-Thought and Code Interpreters
In real-world analytical queries, the reasoning capabilities can be essential, yet LLMs face limitations in this area [24]. To overcome these constraints, advanced methods such as Chain-of-Thought (CoT) [55] and Code-Interpreter (CI) [62] have been introduced. CoT prompts LLMs to generate a series of intermediate reasoning steps, whereas CI lets the LLM produce executable codes and then invoke an external executor to derive the answer. In this section, we evaluate the efficacy of basic generation, the CoT approach, and CI methods using the numerical reasoning dataset FinHybrid. We use the top-5 chunks retrieved with OpenAI’s embedding as the context and benchmark GPT-4-Turbo, Llama-3-8B, and the code-tailored CodeLlama models. The CoT approaches are implemented with step-wise instructive prompts, while the CI method asks LLMs to produce Python codes if necessary (more details in Appendix  B.1).
@@
code 94.4% (no code is generated, directly response) ✗
@@
Remark. As illustrated in Table  9, the Chain-of-Thought (CoT) approach outperforms others across all model configurations (see an example in Table  10). The Llama-3-8B model shows improvements when incorporating codes, yet CoT methods still prove superior. CodeLlama models, however, struggle with generating viable code in this scenario with lengthy context and ambiguous code-gen instructions. GPT-4-turbo exhibits a native ability to produce step-by-step explanations, leading to equivalent performance between basic generation and CoT methods. It is worth noting that while code interpreter has been demonstrated capable in handling numerical and tabular data [26, 65], its efficacy in document analysis is hindered by the unstructured tables and lengthy context.
@@ 5 Conclusion
In this paper, we propose a novel benchmark to assess Retrieval Augmented Generation (RAG) methodologies in real-world document analysis scenarios. Our benchmark features diverse question types and encompasses thousands of unstructured documents with expert labels from financial and other domains; meanwhile, we discussed interesting findings from our initial evaluations. We hope the benchmark will shed light on future research and production in unstructured document analyses.
@@ B.1 Prompt Templates for LLMs
B.2 Settings of Indexing and Retrieval Experiments
In our indexing and retrieval experiments, we first extract raw text from documents, then segment it into 3000-character (about 500 words) chunks using the recursive-split method [45], ensuring a 10% overlap to mitigate information loss. Then the index is constructed on these chunks for the retrieval task. For evaluation, human-annotated factual evidence serves as the ground truth for retrieval. We measure evidence presence by calculating the ratio of the Longest Common Subsequence (LCS) length to the evidence length.
B.3 Long-context Policy

### v2

@@ 1 Introduction
We will actively update the benchmark suite and incorporate more state-of-the-art RAG solutions and LLMs in our benchmark. We hope such efforts will shed light on future research and production of RAG- and LLM-based document analysis.
@@ 3 Dataset: UDA
In this section, we outline the composition and construction of UDA. Each data item within UDA is logically structured as a triplet $(D,q,a)$, where $D$ represents a complete unstructured document, $q$ denotes a question raised from the document, and $a$ signifies the ground truth answer (refer to the data example in Appendix  A). To mirror the authenticity of real-world applications, the documents are retained in their original file formats without parsing or segmentation.
@@
Arithmetic What was the percentage increase in cash dividends from 2015 to 2016? $(0.29-0.25)\div 0.25*100\%=16\%$
Dataset Composition.
@@
We first collect the Q&A labels from the open-released datasets (i.e., source datasets), which are all annotated by human participants. Specifically, our FinHybrid is based on the financial numerical reasoning dataset FINQA [12], which is constructed based on the public earnings reports of S&P 500 companies. TatHybrid is derived from TAT-DQA [66], whose Q&A pairs are accompanied by the document snapshot of 1 to 3 pages from public financial annual reports. Both PaperTab and PaperText are based on Qasper [13], a reading comprehension dataset based on NLP research papers. FetaTab is built upon FetaQA [36], a question-answering dataset for tables from Wikipedia pages. NqText is derived from the widely used Q&A dataset, Natural-Questions [28], which uses the Wikipedia pages as context. Its questions are collected from the Google Search engine, and the answers are human-annotated.
@@
We conduct a series of essential constructing actions after collecting the Q&A labels from the source datasets. The integrity of original documents is crucial for the fidelity of document analysis. However, most of the source datasets only offer well-parsed and segmented partial content without the complete document. To address this problem, we perform a comprehensive source-document identification process, including retrieval, verification, and cleaning. (1) Retrieval: the original documents are sourced from certain platforms, such as Wikipedia [58] and arXiv [2]. We conduct retrieval on these platforms using the metadata from the source datasets, such as titles, timestamps, and document types. (2) Verification: we verify the accuracy of the retrieved document by cross-referencing the content fragments in existing datasets, collecting only exact matches. Specifically, we employ the pypdf [41] library for automatic data parsing and comparison, supplemented by manual inspection. (3) Cleaning: we remove the documents that are inaccessible or not available, such as damaged files and withdrawn papers.
Then we embark on a rigorous matching and reorganization effort, enhancing the data quality and forming complete triplet data pairs, i.e., document-question-answer. This involves the following transformations: (1) data cleaning by removing the Q&A pairs or documents lacking essential answers; (2) standardizing diverse data formats into consistently structured tables and JSON files, addressing the heterogeneity of presentation across different sub-datasets; (3) categorizing the queries in Qasper into table-centric and text-centric, thus forming the PaperTab and PaperText subsets for different application patterns; (4) converting HTML-token-based data type into natural language, forming our user-friendly NqText dataset (5) manually annotating some tables within Qasper to serve as references for evaluating parsing strategies.
@@ 4 Benchmarks and Evaluations
To evaluate the quality of LLM-generated answers, we apply widely accepted span-level F1-score [42] in PaperTab, PaperText, FetaTab, and NqText datasets, where ground-truth answers are in natural language and the source datasets also utilize this metric. We treat the prediction and ground truth as bags of words and calculate the F1-score to measure their overlap. In financial analysis, the assessment becomes more intricate due to numerical values. For the TatHybrid dataset, we adopt the numeracy-focused F1-score, introduced by Zhu, et al. [67], which considers the scale and the plus-minus of numerical values. In the FinHybrid dataset, where answers are always numerical or binary, we rely on the Exact-Match metric but allow for a numerical tolerance of $1\%$, accounting for rounding discrepancies. Furthermore, we also incorporate the LLM-based method for a more comprehensive evaluation (more details in Appendix  B.5). To assess the effectiveness of retrieval strategies, we identify the factual evidence in retrieved chunks using the relative length of the Longest Common Subsequence (LCS) [40] instead of the exact match, because extracted PDF data chunks often include extraneous symbols that can hinder exact matches while still containing crucial evidence.
@@
Following prior works in [4, 60, 68], we focus on zero-shot LLM generation, yet add an extra formatting example to align the output with the desired pattern (refer to Appendix B.1). For the GPT-4 model, we leverage the Azure-OpenAI API to access GPT4-Turbo-1106-Preview with the context window of 128k. Other open-source models are obtained from Huggingface, and we always use the instruct-tuned version. The inference is done on 4 NVIDIA-A100 GPUs. To reduce the compute costs, we randomly sample 1201 documents (in PDF format) accompanied by 2503 question-answer pairs to form our evaluation set (detailed in Table  4). We believe it serves as a practical performance indicator for real-world scenarios.
@@ 4.1 Evaluating Data Parsing
We evaluate various parsing methods to extract tabular information from PDF files and analyze their influence on the downstream tasks, utilizing the table-based questions from PaperTab and FinHybrid (more details in Appendix  B.2). As shown in Figure 2, each question is paired with a PDF page that contains the clue tables; doing so prevents inaccurate retrievals. The tabular data are parsed into text and merged with the rest of the text content as the input context to the LLM.
@@
We evaluate several existing approaches of table parsing: (1) Raw text extraction, which employs a PDF text extractor [41] to extract all the characters. (2) Classic Computer Vision (CV) based approach, which often performs layout detection and OCR extraction at the same time. We follow [55] to use Yolox [17], Tesseract [50] and TableTransformer [51] models together. (3) CV + LLM method, which further employs an LLM to transform the outputs of (2) into Markdown tables. (4) For the advanced multi-modal approach, we employ the latest GPT-4-Omni [37] to convert image-based document tables into Markdown format. (5) The manually-verified well-parsed tables serve as the parsing ground truth.
Table 5: Performance scores (EM or F1) of LLMs using varying parsing strategies on table-based Q&A tasks.
@@
PaperTab (F1) GPT-4-Turbo 42.8 44.3 42.4 38.6 40.7
Llama-3-8B 35.8 37.7 36.5 34.6 32.1
Table 5 reveals that GPT-4-Omni outperforms other approaches, while surprisingly, raw text extraction also yields decent results. We found that the queried tables are relatively simple; the structural markers from the raw text, such as line-breakers and space, are often adequate for LLMs to understand the table. Classic CV methods, if not meticulously tuned, may struggle in handling non-standard table presentations, i.e., edge cases (see an example in Figure  3). Additionally, employing GPT-4-Omni directly for question-answering scores 69.8 and 35.4, lower than sequentially parsing and generating with GPT-4 (i.e., 72.4 and 44.3).
@@
We also observe from the FinHybrid dataset that the GPT-4 model shows a modest 5.7% improvement with well-parsed data over raw-text data, while the much smaller Llama-3-8B offers a significant 15% enhancement, suggesting that compact models with a limited capability of parsing table layouts, may benefit more from enhanced parsing. In the PaperTab dataset, where completely accurate information is less critical, GPT-4-Omni and raw-text parsing could even outperform well-parsed tables by preserving structural cues that highlight important elements for LLM interpretation.
@@
Evaluations here suggest (1) CV-based parsing methods may require adaptation for edge cases before they can be useful; (2) smaller LLMs may be impacted more by uncleaned input data, when requiring accurate and specific information.
@@ 4.2 Indexing and Retrieval
We use the relative length of the Longest Common Subsequence (LCS) to demonstrate the presence of human-annotated evidence in retrieved chunks (more experimental details in Appendix B.3). As shown in Table  6, OpenAI’s text-embedding-3-large model excels in most datasets except for FinHybrid, where the simpler BM-25 approach intriguingly outperforms. This could be attributed to the fact that financial queries often contain more precise details, such as dates or keywords; this aligns well with the direct keyword-matching of BM-25.
@@
We also conduct end-to-end experiments to verify the impact of the retrieval quality. We evaluate the answer quality with top-5 chunks retrieved using OpenAI embeddings or with the human-annotated evidential chunks. As shown in Table 7, providing more relevant context to LLMs improves the answers in most cases, particularly in arithmetic-reasoning tasks such as FinHybrid and TatHybrid. Interestingly, for the FinHybrid dataset, the Llama-3-8B model achieved a score of 51.0 when given accurate context, outperforming GPT-4 with model-retrieved chunks. However, for knowledge-based questions from the NqText and FetaTab, the answer quality remains less affected. This is because complex numerical reasoning demands more precise evidence for accurate arithmetic operations, whereas LLMs can leverage a wide range of narrative information to derive answers to knowledge-based questions.
@@
We compare RAG-based methods with long-context LLMs, utilizing GPT-4-Turbo with a 128k context window and Qwen-1.5-7B with a 32k context window. Due to the high cost of long context inference, we conduct this experiment on a subset of 600 documents (more details in Appendix  B.4). The results are demonstrated in Table  8.
@@
Remark. We notice that for free-form or knowledge-based tasks (i.e., paper-based and wiki-based Q&A), RAG and long-context solutions demonstrate comparable capability. Conversely, in tasks with more numerical reasoning (i.e., financial Q&A), long-context LLMs fail to match the performance of RAG. In such use cases, the excess of verbose content might hinder long-context LLMs in pinpointing facts and performing numerical reasoning effectively (see an example in Table  10). Additionally, RAG is likely to surpass the long-context mechanism more in the smaller LLM, given its constrained capacity to handle large volumes of data.
@@ 4.4 Evaluating Chain-of-Thought and Code Interpreters
In real-world analytical queries, reasoning capabilities can be essential, yet LLMs face limitations in this area [25]. To overcome these constraints, advanced methods such as Chain-of-Thought (CoT) [57] and Code-Interpreter (CI) [65] have been introduced. CoT prompts LLMs to generate a series of intermediate reasoning steps, whereas CI lets the LLM produce executable codes and then invoke an external executor to derive the answer. In this section, we evaluate the efficacy of basic generation, the CoT approach, and CI methods using the numerical reasoning dataset FinHybrid. We use the top-5 chunks retrieved with OpenAI’s embedding as the context and benchmark GPT-4-Turbo, Llama-3-8B, and the code-tailored CodeLlama models. The CoT approaches are implemented with step-wise instructive prompts, while the basic CI method asks LLMs to produce Python codes if necessary (more details in Appendix  B.1).
@@
code 94.4% (no code is generated, direct response) ✗
@@
Remark. As illustrated in Table  9, the Chain-of-Thought (CoT) approach outperforms others across all model configurations (see an example in Table  10). The Llama-3-8B model shows improvements when incorporating codes, yet CoT methods still prove superior. CodeLlama models, however, struggle with generating viable code in this scenario with lengthy context and ambiguous code-gen instructions. GPT-4-turbo exhibits a native ability to produce step-by-step explanations, leading to equivalent performance between basic generation and CoT methods. It is worth noting that while code interpreter has been demonstrated capable of handling numerical and tabular data [27, 68], its efficacy in document analysis is hindered by the unstructured tables and lengthy context, if just using the basic code strategy.
@@ 4.5 End-to-End Evaluations
Llama-3-8B-NoRAG 18.0 3.6 6.0 13.0 16.2 47.7 21.7
GPT-4-turbo-NoRAG 17.6 0.4 3.0 15.3 18.9 48.6 19.6
Table 12: Performance scores with and without RAG
LLM and Strategy FinHybrid TatHybrid PaperTab PaperText FetaTab NqText
GPT-4 45.9 43.5 40.3 45.8 61.5 37.4
GPT-4-NoRAG 0.4 3.0 15.3 18.9 48.6 19.6
Llama-3-8B 37.9 22.5 35.5 42.3 56.6 31.7
Llama-3-8B-NoRAG 3.6 6.0 13.0 16.2 47.7 21.7
@@
Additionally, Table  12 presents the results when the LLMs respond to the questions with versus without RAG support. A significant decline in accuracy is observed when RAG is absent, underscoring the LLMs’ deficiency of related internal knowledge and their reliance on external documents. This reinforces the effectiveness of our benchmark for the evaluation and exploration of RAG approaches.
@@ 5 Conclusion
In this paper, we propose a novel benchmark to assess Retrieval Augmented Generation (RAG) methodologies in real-world document analysis scenarios. Our benchmark features diverse question types and encompasses thousands of unstructured documents with expert labels from financial and other domains. Meanwhile, we discuss interesting findings from our evaluations, covering data parsing, information retrieval, long context mechanism, generating strategies and end-to-end performance. We believe our benchmark will advance future research and production in unstructured document analysis.
6 Limitations
Despite the valuable contributions of this study, we acknowledge its limitations: (1) While the efficacy of parsing strategies was evaluated through the downstream Q&A performance, a direct comparison of the parsed content was not undertaken. This stems from the absence of well-defined standards for direct quality assessment in the scenario of document understanding. (2) This study did not extend to the in-depth analyses of noise sensitivity and hallucination. We will delve into these topics and conduct detailed analyses in our future work.
@@ B.1 Prompt Templates for LLMs
B.2 Settings of Parsing Experiments
In our parsing experiments, we utilize 341 table-based questions from FinHybrid and 100 questions from PaperTab. FinHybrid provides cleanly extracted tables which we formatted into Markdown text as our well-parsed ground truth. To get the well-parsed content in PaperTab, we first use GPT-4-Omni to convert table images to Markdown, followed by manual corrections for inaccuracies. Since Markdown doesn’t support the nested structure, we replicate the root content to represent hierarchical relationships.
B.3 Settings of Indexing and Retrieval Experiments
In our indexing and retrieval experiments, we first extract raw text from documents, then segment it into 3000-character (about 500 words) chunks using the recursive-split method [47], ensuring a 10% overlap to mitigate information loss. Then the index is constructed on these chunks for the retrieval task. For evaluation, human-annotated factual evidence serves as the ground truth for retrieval. We measure evidence presence by calculating the ratio of the Longest Common Subsequence (LCS) length to the full evidence length.
B.4 Long-context Policy